OpenStudy (anonymous):

How many different strings can be made from the letters in MISSISSIPPI, using all the letters and having no consecutive S's?

OpenStudy (anonymous):

the answer is c(8,4) x 7!/(4!2!) I don't get where they get 4!2! from

OpenStudy (anonymous):

\[\frac{11!}{4!4!2!}\]

OpenStudy (anonymous):

why

OpenStudy (anonymous):

11 letters and 4S, 4I, 2P

OpenStudy (zarkon):

mississippi treat the m,i's,p's as dividers MIIIPPI we will put the S's in between the bars...there are \(_8C_4\) ways to do this then we need to permute the 'dividers'. (use the multinomial coefficient) there are 7 dividers...1 single letter and one group of 2 and one group of 4 this gives \(\displaystyle \frac{7!}{1!2!4!}=\frac{7!}{4!2!}\)